A pendulum clock is taken 1 km inside the earth from mean sea level. Then the pendulum clock
  • Loses 13.5 sec per day
  • Gains 13.5 sec per day
  • Loses 7 sec per day
  • Gains 7 sec per day​​​​
Pls provide the answer with appropriate method and along with calculation and also explain step by step ...pls 

Grade:11

1 Answers

Vikas TU
7 years ago
At sea level,
gravitational constant that is
g = GM/R
and Time period,
T = 2pi*root(l/g)
at below 1 km,
g’ = GM/(R-1000)
T’ = 2pi*root(l/g’)
T/T’ = root(g’/g) = > root(R/(R-1000))
T’ = T/ root(R/(R-1)) => T/root(6400/6399) = > T/root(1.000156274) => T/1.000078134
or
T’ = 0.999921872*T
putting value of T for appropriaate T we get
T’ approx 13.6 seconds.